What is the slope and y-intercept of 5y+4x=5
Answer by AnlytcPhil(1116)

What is the slope and y-intercept of 5y+4x=5

Rule: 

1. Solve for y.

2. The slope equals the coefficient of x on the right side.

3. The y-intercept is the constant term on the right side

                      5y + 4x = 5

                           5y = 5 - 4x

Divide every term by 5

                           5y     5     4x
                          ———— = ——— - ————
                            5     5      5

                           1      1
                           5y     5     4x
                          ———— = ——— - ————
                            5     5      5         
                            1     1
                                      4x
                             y = 1 - ————
                                       5

                                      4
                             y = 1 - ———·x
                                      5

Now that we have solved for y, the coefficient of x on 
the right side is -4/5 so that's the slope. The constant
term on the right side in 1, so that's the y-intercept.
Actually the y-intercept is the point (0,1) on the y-axis.

Edwin
AnlytcPhil@aol.com